find exact value
cos (-15 degrees)

Answer by lwsshak3(11628)

find exact value
cos (-15 degrees)=cos(45-60)
=cos45cos60+sin45+sin60=√2/2*1/2+√2/2*√3/2=√2/4+√6/4=(√2+√6)/4
Check:
cos(-15˚)≈0.9659
Exact value as computed=(√2+√6)/4≈0.9659